OpenStudy (babyslapmafro):

Please help me with a part of the following problem. I am asked to find parametric equations for the line through P1 and P2 and also for the line segment joining those points. P1(3,-2) P2(5,1) I found the parametric equations for the line through P1 and P2 to be: x=2t+3 y=3t-2 but I do not know how to find a parametric equation for the line segment joining those points.

OpenStudy (badhi):

I think its the same answer you have written but with, t being bounded. So that the line will not go further than the point P1 and P2

OpenStudy (anonymous):

is the parametric equations that u found in first place right?

OpenStudy (babyslapmafro):

no, i just realized i typed it wrong

OpenStudy (babyslapmafro):

ok i changed it

OpenStudy (anonymous):

we must bound t

OpenStudy (loser66):

the parametric equation for the line segment from P1 to P2 is x = 3+2t y = -2 +2t I agree with BAdhi, they are same.

OpenStudy (babyslapmafro):

I'm still not sure how to 'bound t'

OpenStudy (babyslapmafro):

How do I determine which values bound t?

OpenStudy (phi):

starting with your solution x=2t+3 y=3t-2 when t=0 you get x=3, y= -2 (3,-2) = P1 when t= 1 you get x=5, y=1 (5,1) = P2 so I would limit t: 0 ≤ t ≤ 1 if you only want the line segment between P1 and P2 inclusive